Timeline


and

09/12/18:

14:15 ServerShutdownProcedure edited by jennifer
(diff)
14:14 ServerShutdownProcedure edited by jennifer
(diff)
13:49 ServerShutdownProcedure created by ed
13:31 WikiStart edited by ed
(diff)
13:16 Ticket #357 (Network Outage at Goddard 9/12/18) closed by jennifer
fixed

09/11/18:

12:12 Ticket #357 (Network Outage at Goddard 9/12/18) created by alisdair
Goddard is having a planned network outage 09/12/18 starting at 11:15 AM …

09/10/18:

15:19 Ticket #352 (Ticket emails to Jacob and Joe G. are not being delivered) closed by alisdair
fixed: Update was done by Joe G. Ticket system and virtualsolar.org emails seem to be working well now. Closing this ticket.
15:17 Ticket #355 (SPF record change for virtualsolar.org) closed by alisdair
fixed: Gurman made the appropriate changes and Jacob just tested the tickets alias out for me, as I will here. It all seems to be working so will close this ticket.

09/07/18:

16:06 Ticket #355 (SPF record change for virtualsolar.org) reopened by alisdair

09/06/18:

15:01 Ticket #355 (SPF record change for virtualsolar.org) closed by alisdair
fixed: New SPF record is live.
14:56 Ticket #356 (KSO data provider now supports SAMPLE.) closed by alisdair
fixed
14:56 Ticket #356 (KSO data provider now supports SAMPLE.) created by alisdair
Added support for SAMPLE to the KSO data provider. Code is in place at …

09/05/18:

14:58 Ticket #355 (SPF record change for virtualsolar.org) created by alisdair
Asked Joe G. to update SPF record for virtualsolar.org so I can test …
09:14 Ticket #354 (Returned results are not date sorted. Add /sort flag.) created by alisdair
When doing a search the order that the results are returned in is not …

09/04/18:

15:32 Various edited by alisdair
(diff)
15:32 SortingResults edited by alisdair
(diff)
15:31 SortingResults edited by alisdair
(diff)
14:06 twRAID edited by niles
(diff)
13:51 Various edited by alisdair
(diff)
13:50 Various edited by alisdair
(diff)
13:49 SortingResults created by alisdair
13:39 Various edited by alisdair
(diff)
13:38 Various edited by alisdair
(diff)
13:38 Various edited by alisdair
(diff)
08:34 Ticket #353 (VSO1 /opt raid disk replacement) closed by jacob
fixed
08:34 Ticket #353 (VSO1 /opt raid disk replacement) created by jacob
On Friday of last week the sda2 raid went down again, taking with it …

08/26/18:

15:08 Ticket #352 (Ticket emails to Jacob and Joe G. are not being delivered) created by alisdair
For some reason ticket emails to Jacob and Joe G. are not being delivered. …
15:08 Ticket #346 (SDO downloads seem impeded) closed by niles
fixed: I'm resolving this as the 504 status returns have gone now that both NSO and presumably SDAC are up to date. It needs to be watched, though, if we find that we are again not keeping up we need to revisit.

08/24/18:

14:15 Ticket #351 (Problem downloading AIA data from SDO6 (SDAC) using IDL client) created by alisdair
[…]
11:20 Ticket #350 (Configure software from vso03 on netdrms02) created by niles
Need to get the NSO VSO instance and the data providers going on …
11:17 Ticket #349 (Remote SUMS on netdrms02) created by niles
Having established ssh access to the JSOC from netdrms02 (Thanks Art!) I …

08/23/18:

15:24 Ticket #348 (Issues with vso03) created by niles
Recently vso03 has been halting in a catastrophic way. Twice it got into a …
15:01 Ticket #347 (We need more backup disks) created by niles
We need more cold spares. I (Niles) need to order both an INTEL 200GB and …
14:58 Ticket #346 (SDO downloads seem impeded) created by niles
Executive summary : I think downloads from the JSOC at NSO are not …

08/21/18:

11:16 Ticket #341 (VSO03 @ NSO down) closed by alisdair
fixed
10:55 Ticket #345 (sdo5 (SDAC) via and IDL query returns wrong path for drms_export CGI at ...) closed by alisdair
fixed: Deleted the /var/www/cgi-bin/drms_test directory on netdrms01.nispdc.nso.edu as it was no longer necessary after the permanent fix. Closing this ticket.

08/20/18:

15:58 Ticket #345 (sdo5 (SDAC) via and IDL query returns wrong path for drms_export CGI at ...) created by alisdair
The problem is that SDO5 (at SDAC) is responding with the wrong path for …

08/17/18:

12:31 Ticket #344 (Monitoring script Changes) closed by jacob
fixed
12:31 Ticket #344 (Monitoring script Changes) created by jacob
Following the punctual and descriptive, but perhaps a little too frequent …

08/16/18:

12:18 Ticket #343 (VSO1 /opt raid down) closed by jacob
fixed: Thanks in no part to my efforts, the databases are back up on vso1.
11:45 Ticket #343 (VSO1 /opt raid down) created by jacob
The RAID housing most of the databases for SDAC data is down at the …
11:22 Ticket #288 (SDO6 testing) closed by ed
fixed
Note: See TracTimeline for information about the timeline view.